Earleton Manufacturing Company has $2 billion in sales and $500,000,000 in fixed assets. Currently, the company's fixed assets are operating at 85% of capacity.

  1. What level of sales could Earleton have obtained if it had been operating at full capacity? Write out your answer completely. Round your answer to the nearest cent.

    $   

  2. What is Earleton's target fixed assets/sales ratio? Round your answer to two decimal places.

    %

  3. If Earleton's sales increase 20%, how large of an increase in fixed assets will the company need to meet its target fixed assets/sales ratio? Write out your answer completely. Do not round intermediate calculations. Round your answer to the nearest dollar.

    $  
Answer #1

a). Full Capacity sales = Actual Sales / % of capacity at which Fixed Assets are operated

= $2,000,000,000 / 0.85 = $2,352,941,176

b). Target Fixed Assets/Sales Ratio = Fixed Assets / Full Capacity Sales

= $500,000,000 / $2,352,941,176 = 0.2125, or 21.25%

c). New Sales = Current Sales * (1 + g) = $2,000,000,000 * (1 + 0.20) = $2,400,000,000

Change in Fixed Assets = Target Fixed Assets/Sales Ratio * Change in Sales

= 0.2125 * [$2,400,000,000 - $2,000,000,000]

= 0.2125 * $400,000,000 = $85,000,000